# CLASSMix: Adaptive stain separation-based contrastive learning with pseudo labeling for histopathological image classification

### v1

@@ Abstract
Histopathological image classification is one of the critical aspects in medical image analysis. Due to the high expense associated with the labeled data in model training, semi-supervised learning methods have been proposed to alleviate the need of extensively labeled datasets. In this work, we propose a model for semi-supervised classification tasks on digital histopathological Hematoxylin and Eosin (H&E) images. We call the new model Contrastive Learning with Adaptive Stain Separation and MixUp (CLASSMix). Our model is formed by two main parts: contrastive learning between adaptively stain separated Hematoxylin images and Eosin images, and pseudo labeling using MixUp. We compare our model with other state-of-the-art models on clear cell renal cell carcinoma (ccRCC) datasets from our institution and The Cancer Genome Atlas Program (TCGA). We demonstrate that our CLASSMix model has the best performance on both datasets. The contributions of different parts in our model are also analyzed.
@@
Contrastive learningStain separationPseudo labelingMixUpDigital histopathological imagesSemi-supervised learning
@@ 1 Introduction
Digital histopathological image analysis plays a crucial role in disease diagnosis and treatment optimization. The most commonly used histopathological images are Hematoxylin & Eosin (H&E) stained whole slide images (WSIs) that help the differentiation of various tissue sample features [15]. H&E staining involves the use of two dyes, Hematoxylin and Eosin, which selectively stain different components of the tissue sample. Hematoxylin stains the acidic components of the tissue sample, such as the cell nuclei, and has a blue-purple color, while Eosin stains the basic components of the tissue sample, such as the cytoplasm and extracellular matrix, and exhibits a pink color. In recent years, many efforts have been made to the automatic analysis of images.
@@
With the rapid development of computing power and advanced algorithms, deep learning has been widely used in the field of digital histopathological image analysis for automating disease diagnosis and performing auxiliary image analysis [8, 36, 43, 45]. However, achieving a highly accurate model requires a substantial amount of training labels, demanding significant time and effort from human experts. Various strategies have been proposed to mitigate this challenge. Weakly-supervised learning become the accepted method in histopathology because of the easy availability of case-level labels from medical reports. Self-supervised learning is also commonly used to train the backbone. Nevertheless, semi-supervised learning remains a relatively less explored area in histopathology.
@@
One area worthy of further investigation with deep learning models is kidney cancer. Kidney cancer ranks among the most prevalent cancers globally. It is estimated that around 76,080 new cases were diagnosed with cancers of the kidney and renal pelvis and there are around 13,780 deaths resulted from it in the US in 2021 [38]. Clear cell renal cell carcinoma (ccRCC) stands out as the predominant subtype that dominates the kidney cancer cases [4]. Consequently, further research on cancer detection and classification using ccRCC images holds paramount importance for disease diagnosis and early patient treatment. In this paper, we contribute new annotated Utah ccRCC dataset and The Cancer Genome Atlas Program (TCGA) ccRCC dataset. The Utah ccRCC dataset consists of 49 WSIs, while TCGA ccRCC dataset comprises 420 WSIs. Even though multi-instance learning, requiring only slide-level annotations, is widely used in the histopathology field [32, 35, 42, 50], it requires a larger number of WSIs for effective training. Moreover, slide-level multi-instance learning has some prerequisites, like certain classes should have both positive and negative cases, which can’t always be fulfilled in our ccRCC datasets. Therefore, we decided to use polygons to mark regions in a subset of WSIs and assign labels to those polygons. The patches can be cropped inside those annotated polygons to collect labeled samples for performing patch-level classification tasks. To fully utilize the WSIs, we collect the patches outside annotated polygons and from WSIs that are not annotated at all to gather unlabeled samples. Consequently, we form a semi-supervised classification task based on the labeled and unlabeled patches. Notably, a patch-level classification model serves as a precise tool to expedite pathologists’ identification of cancerous regions within WSIs.
@@
In our work, we propose a new semi-supervised model for histopathological image classification: Contrastive Learning with Adaptive Stain Separation and MixUp (CLASSMix). The main portions of our model are contrastive learning and pseudo labeling. A simple but effective margin loss is built for contrastive learning between adaptive stain separated Hematoxylin images and Eosin images. The pseudo labeling with MixUp is also applied to provide more samples and pseudo labels in classification training. The two methods focus on different aspects to fully utilize the information in both labeled and unlabeled samples. Our new model is an extension of [54] presented in MICCAI 2022 workshop with substantial modifications and great improvements on classification results. Instead of using a globally fixed stain separation matrix in Optical Density space, a simple and efficient slide-by-slide stain separation from [37] is applied to utilize domain knowledge of color patterns across slides. We not only perform image augmentations on Hematoxylin images and Eosin images, but also apply augmentations on original RGB images before stain separation process to strengthen the augmentation process. Additionally, considering the benefit of pseudo labeling in famous semi-supervised learning methods [47] [3], pseudo labeling with MixUp [56] is adopted to mix samples and then create mixed one-hot encoding labels. In this paper, we evaluated multiple state-of-the-art semi-supervised and self-supervised learning models on Utah ccRCC dataset and TCGA ccRCC dataset. Our CLASSMix model demonstrates the best classification results. We further did ablation studies to carefully analyze the contributions of different parts in our model.
@@
  • We proposed our new CLASSMix model for semi-supervised learning. We applied contrastive learning on Hematoxylin images and Eosin images after performing slide-level stain separation, and pseudo labeling with MixUp to generate more mixed samples and pseudo labels.
  • We provide new Utah ccRCC dataset from our institution and TCGA ccRCC dataset that contain tile-level labels on 49 WSIs and 150 WSIs separately, with extra 270 unlabeled WSIs from TCGA. The Utah ccRCC dataset and its annotations are publicly available upon transfer agreement. The annotations on TCGA ccRCC dataset are publicly available.
  • We tested various state-of-the-art semi-supervised learning and self-supervised learning methods on ccRCC datasets and benchmarked their performances. The results of CLASSMix always have the best performance among state-of-the-art models.
@@ 2.1 Weakly-supervised learning
Weakly-supervised learning [28, 39, 58] leverages weak labels to help guide training process. A typical case of weakly-supervised learning is multiple instance learning (MIL) [16, 24]. Various MIL approaches, such as instance-based and embedding-based methods, have been proposed. The instance-based approach [25, 13] involves training a classifier using bag labels at the instance level, then aggregating the predicted instance labels to form bag label predictions. However, due to the noise in instance labels, the performance of this aggregation may be impacted. On the other hand, the embedding-based approach [19, 34] first creates a bag representation from individual instance representations, followed by training a classifier on these bag representations. Research has shown that embedding-based approaches are generally more effective than instance-based methods. The attention-based MIL method, proposed in [26], suggests that instead of treating all patches within a bag equally, assigning varying importance scores to patches, particularly the more discriminative ones, is more effective. This method involves computing attention scores from instance representations to reflect the importance of respective patches and forming the bag representation through a weighted average of these instances. [51] proposed a hierarchical attention-guided MIL that effectively identifies important areas at different scales in WSIs. This method combines several attention techniques to form a comprehensive group representation. Furthermore, the recent Self-ViT-MIL approach by [18], which combines self-supervised learning with Vision Transformers (ViTs) [14] and MIL, has shown promising results, even in comparison to fully-supervised methods.
In digital histopathological image applications, MIL can be applied when only whole slide image-level labels are provided, but the specific regions that contribute to the labels are not given. Despite the convenience of annotation work, weakly-supervised learning can’t deal with small datasets so well, as it needs adequate amount of slides to have enough labels. Additionally, weakly-supervised learning struggles with multiple classes which almost always co-exist in slides or a certain class that lacks positive or negative labels in datasets. For example, in our Utah ccRCC dataset, cancer is consistently present in all WSIs, resulting in a lack of negative labels for the cancer category in the context of MIL.
@@ 2.2 Self-supervised learning
Self-supervised learning (SSL) [27, 30] learns useful patterns from data itself without explicit labels provided by humans. Self-supervised learning allows models to first pre-train on a large unlabeled dataset, where effective feature representations can be learned. Examples include image inpainting, predicting rotations, or colorizing images. Then the final layers of the models can be optimized in specific downstream tasks, reducing the need of extensively labeled data. SimCLR [9, 10] presents a seminal contrastive learning to train feature representation, where augmented views of an image (positive samples) are minimized within that image while being maximized against augmented views of other images (negative samples). MoCo [11, 12, 21, 22, 20] strengthens SSL with a dynamic dictionary, tapping into past batches for more negative samples, and a slowly updating momentum encoder for stable global representations. In contrast, Barlow Twins [53] omits negatives, focusing on refining features by measuring cross-correlation between positive samples. BYOL [17] and DINO [7] utilize a teacher-student setup, where the student refines its understanding by predicting the teacher’s representation, while teacher network is updated through the slow-moving average of student’s parameters. Clustering-based methods like Deep Clustering [5] and SwAV [6] define pseudo-labels, sorting images into clusters, yielding great performance on downstream tasks. Mingu et al. [29] benchmarked SSL methods on pathology datasets, showcasing consistent enhancements in histopathology tasks. In CS-CO [52], the encoders for Hematoxylin images and Eosin images are also adopted. Instead of contrastive loss between Hematoxylin images and Eosin images introduced in our work, CS-CO calculates contrastive loss between different augmentations. In addition, CS-CO is a self-supervised learning method that requires encoders to generate a visual representation containing enough information to recover images in cross-stain prediction, while our semi-supervised CLASSMix only pursues shared latent features containing enough information to perform classification tasks. Self-supervised learning is especially helpful when there is a lack of labeled data for training. While self-supervised learning benefits from unlabled data, semi-supervised learning has the advantages that allows both unlabeled data and labeled data to be trained at the same time and use the knowledge learned from labeled data to better utilize unlabeled data.
@@ 2.3 Semi-supervised learning
Semi-supervised learning [40, 48] involves training a model with both labeled and unlabeled data to enhance training outcomes, which reduces reliance on a limited number of annotated samples. Semi-supervised learning has superiority especially when obtaining fully labeled data is prohibitively expensive or impractical. In the realm of digital histopathological image analysis, getting a large amount of labeled data requires great effort from well-trained experts, which is expensive and time consuming. However, it is much easier to acquire unlabeled data such as unannotated WSIs. One type of semi-supervised learning is consistency regularization which seeks agreement among the same input with different views, augmentations or epochs. For example, Temporal ensembling [31] aims to reach consensus in prediction of labels between current epochs and previous epochs. [44] tries to make classifier’s prediction consistent across various transformations. More up-to-date consistency regularization methods, such as FixMatch [47] and MixMatch [3], have been developed. FixMatch assigns pseudo labels to unlabeled samples when the predicting confidence is high with weak augmentation of input data, then the unlabeled data will be treated as labeled data and trained after strong augmentation. MixMatch is another consistency regularization method where pairs of labeled/unlabeled input data are linearly combined to create "Mixup" data. The label of new created data is a weighted average of the original labels followed by a sharpening process. Inspired by self-supervised learning, contrastive learning idea is also integrated in semi-supervised learning [1, 46, 54]. Instead of seeking agreements in label predictions, contrastive learning focuses on forming positive pairs and negative pairs. Positive pairs originate from the same inputs with different views or augmentations, while negative pairs involve different inputs. The method encourages a shorter distance between positive pairs and a longer distance between negative pairs in the feature representation space.
Figure 1: The block-diagram of our Contrastive learning with adaptive stain separation and mixup (CLASSMix) for semi-supervised histopathological image classification. Orange part shows adaptive stain separation, where OD means Optical Density space. Green part shows mixup on both labeled and unlabeled samples.
@@ 3 Methods
3.1 An overview of CLASSMix
Figure 1 shows the workflow of our Contrastive learning with adaptive stain separation and mixup (CLASSMix). As shown in orange part of Figure 1, slide-level stain separation is applied for generating Hematoxylin images and Eosin images from original RGB images. The original RGB images, Hematoxylin images and Eosin images are all augmented during training to improve the robustness of CLASSMix model. The model takes Hematoxylin images and Eosin images as input pairs to form different views of input data. The purple and pink boxes in Figure 1 illustrate the H and E ResNet [23] encoders that have the same architecture but separate parameters to generate latent features $f_{H}$ and $f_{E}$. A contrastive loss is proposed for shared latent feature space between H channel and E channel to increase the similarity between $f_{H}$ and $f_{E}$. We take the average of features $f_{H}$ and $f_{E}$ and pass it to a linear+softmax layer to predict the labels. For labeled samples in training, the cross-entropy loss is introduced to measure the difference between predicitons and true labels. Moreover, the main challenge in semi-supervised learning in general is the limited number of labeled samples. Pseudo labeling on unlabeled samples has been widely used in many state-of-the-art semi-supervised learning methods [3, 47, 55]. In MixMatch [3], after pseudo labeling, the model further used MixUp [56] to introduce virtual samples by linear interpolation of two random samples. Inspired by those ideas and to fully utilize unlabeled samples, we first provide pseudo labels to unlabeled samples, then add MixUp on both labeled and unlabeled samples to create virtual samples. The labels of the generated samples after MixUp are set to weighted averages of original labels/pseudo labels with a sharpening process. The green part in Figure 1 shows the workflow of this process for pseudo labeling. We introduce different parts of CLASSMix model in details in the following subsections.

### v4

# CLASS-M: Adaptive stain separation-based contrastive learning with pseudo-labeling for histopathological image classification

@@ Abstract
Histopathological image classification is an important task in medical image analysis. Recent approaches generally rely on weakly supervised learning due to the ease of acquiring case-level labels from pathology reports. However, patch-level classification is preferable in applications where only a limited number of cases are available or when local prediction accuracy is critical. On the other hand, acquiring extensive datasets with localized labels for training is not feasible. In this paper, we propose a semi-supervised patch-level histopathological image classification model, named CLASS-M, that does not require extensively labeled datasets. CLASS-M is formed by two main parts: a contrastive learning module that uses separated Hematoxylin and Eosin images generated through an adaptive stain separation process, and a module with pseudo-labels using MixUp. We compare our model with other state-of-the-art models on two clear cell renal cell carcinoma datasets. We demonstrate that our CLASS-M model has the best performance on both datasets. Our code is available at github.com/BzhangURU/Paper_CLASS-M/tree/main
@@
Contrastive learningStain separationpseudo-labelingSemi-supervised learningDigital histopathological images
@@ 1 Introduction
Digital histopathological image analysis plays a crucial role in disease diagnosis and treatment optimization. The most commonly used histopathological images are Hematoxylin & Eosin (H&E) stained whole slide images (WSIs) that help the differentiation of various tissue sample features  [18]. H&E staining involves the use of two dyes, Hematoxylin (H) and Eosin (E), which selectively stain different components of the tissue samples. Hematoxylin stains the acidic components of the tissue samples, such as the cell nuclei, and has a blue-purple color, while Eosin stains the basic components of the tissue samples, such as the cytoplasm and extracellular matrix, and exhibits a pink color. In recent years, many efforts have been made towards the automatic analysis of H&E images.
@@
With the rapid development of computing power and advanced algorithms, deep learning has been widely used in the field of digital histopathological image analysis for automating disease diagnosis and performing auxiliary image analysis  [9, 39, 46, 48, 53]. However, achieving a highly accurate model requires a substantial amount of training labels, demanding significant time and effort from human experts. Various strategies have been proposed to mitigate this challenge. Multiple instance learning (MIL), a specific type of weakly-supervised learning, has become the accepted approach in histopathology because of the easy availability of case-level labels from medical reports. Self-supervised learning is also commonly used to train the backbone feature extractor network. Nevertheless, semi-supervised learning, despite its success in computer vision, remains a relatively less explored area in histopathology.
@@
Another direction worthy of further investigation with deep learning models is kidney cancer. Kidney cancer ranks among the most prevalent cancers globally. It is estimated that around 76,080 new cases were diagnosed with cancers of the kidney and renal pelvis and there are around 13,780 deaths resulting from it in the US in 2021  [41]. Clear cell renal cell carcinoma (ccRCC) stands out as the predominant subtype that dominates the kidney cancer cases  [4]. Consequently, further research on cancer detection and classification using ccRCC images holds paramount importance for disease diagnosis and early patient treatment. In this paper, we experiment with two ccRCC datasets. Even though MIL approaches, requiring only slide-level annotations, are widely used in the histopathology field  [35, 38, 45, 54], they require a large number of WSIs for effective training. For instance, the Utah ccRCC dataset we experiment with consists of only 49 WSIs. Furthermore, if multiple classes always co-occur in WSIs, MIL approaches can not differentiate between them. This is another issue with the Utah ccRCC dataset which contains only cancer cases; therefore, normal regions always co-occur with cancer regions in WSIs. Finally, MIL methods typically classify WSIs based on only the most discriminative patches leading to poor patch-level accuracy despite demonstrating good slide-level accuracy. This can be a problem even for larger datasets such as The Cancer Genome Atlas Program (TCGA) ccRCC dataset we experiment with which comprises 420 WSIs.
In our approach, pathologists draw approximate polygons at a low resolution to mark regions in a subset of WSIs and assign labels to those regions. We crop patches from inside these annotated regions to collect labeled samples for performing patch-level classification tasks. To fully utilize the WSIs, we also collect patches from outside the annotated regions and from WSIs that are not annotated to gather unlabeled samples. Consequently, we form a semi-supervised classification task based on the labeled and unlabeled patches. Notably, a patch-level classification model serves as a precise tool to expedite pathologists’ identification of cancerous regions within WSIs.
@@
In our work, we propose a new semi-supervised model for histopathological image classification: Contrastive Learning with Adaptive Stain Separation and MixUp (CLASS-M), where CLASS-M can also be understood as a model for CLASSifying Medical images. CLASS-M is an extension of  [58] presented in MICCAI 2022 Workshop on Medical Image Learning with Limited and Noisy Data with substantial modifications, a more thorough experimental analysis and significant improvements in classification accuracy. A simple and effective novel unsupervised loss is formulated for contrastive learning between adaptively stain-separated Hematoxylin images and Eosin images. Additionally, considering the benefit of pseudo-labeling in state-of-the-art general semi-supervised learning methods  [50]  [3], pseudo-labeling with MixUp  [60] is adopted to provide further regularization via additional samples and pseudo-labels in training. The two methods benefit from different regularization effects, contrastive learning with multiple views vs. data augmentation, and hence can be combined to better utilize the information in both labeled and unlabeled samples and improve classification accuracy. Furthermore, instead of using a globally fixed stain separation matrix in Optical Density space as in  [58], a simple and efficient slide-by-slide stain separation based on  [40] is applied to adapt to properties of individual WSIs. We not only perform image augmentations on Hematoxylin images and Eosin images, but also apply augmentations on original RGB images before the stain separation to strengthen the augmentation process. Finally, in this paper, we compare our method to multiple state-of-the-art semi-supervised and self-supervised learning models on the Utah ccRCC dataset and TCGA ccRCC dataset. We also conduct ablation studies to carefully analyze the contributions of different parts in our model.
@@
  • We propose the CLASS-M model for semi-supervised learning. We apply our novel contrastive learning on Hematoxylin images and Eosin images after performing slide-level stain separation. We also use pseudo-labeling with MixUp to further improve classification accuracy.
  • We provide the new Utah ccRCC dataset which has 49 WSIs with patch-level labels, and patch-level labels for 150 WSIs from the TCGA ccRCC dataset. The Utah ccRCC dataset and its annotations will be made available upon transfer agreement. The annotations on TCGA ccRCC dataset are made publicly available.
  • We test various state-of-the-art semi-supervised learning and self-supervised learning methods on both ccRCC datasets and demonstrate that CLASS-M outperforms other state-of-the-art models.
  • The code for our paper is available at github.com/BzhangURU/Paper_CLASS-M/tree/main
@@ 2.1 Weakly-supervised learning
Weakly-supervised learning  [31, 42, 62] leverages slide or case-level labels to help guide the training process. A typical case of weakly-supervised learning is MIL  [19, 27]. Various MIL approaches, such as instance-based and embedding-based methods, have been proposed. The instance-based approach [28, 16] involves modeling a classifier at the instance level, then aggregating the predicted instance labels to form bag label predictions. However, due to the noise in instance labels, the performance of this aggregation may be impacted. On the other hand, the embedding-based approach [22, 37] first creates a bag representation from individual instance representations, followed by training a classifier on these bag representations. Research has shown that embedding-based approaches are generally more effective than instance-based methods. The attention-based MIL method, proposed in  [29], suggests that instead of treating all patches within a bag equally, assigning varying importance scores to patches, particularly the more discriminative ones, is more effective. This method involves computing attention scores from instance representations to reflect the importance of respective patches and forming the bag representation through a weighted average of these instances.  [55] proposed a hierarchical attention-guided MIL that effectively identifies important areas at different scales in WSIs. This method combines several attention techniques to form a comprehensive group representation. Furthermore, the recent Self-ViT-MIL approach by  [21], which combines self-supervised learning with Vision Transformers (ViTs)  [17] and MIL, has shown promising results, even in comparison to fully-supervised methods.
In digital histopathological image applications, MIL can be applied when only whole slide image-level labels are provided, but the specific regions that contribute to the labels are not given. Despite the convenience of annotation work, weakly-supervised learning can’t deal with small datasets so well, as it needs an adequate amount of slides for learning. Additionally, weakly-supervised learning struggles with multiple classes which almost always co-exist in slides or a certain class that lacks positive or negative labels in datasets. For example, in our Utah ccRCC dataset, cancer is consistently present in all WSIs, resulting in a lack of negative labels for the cancer category in the context of MIL.
@@ 2.2 Self-supervised learning
Self-supervised learning (SSL)  [30, 33] extracts useful patterns from data itself without explicit labels provided by humans. Self-supervised learning allows models to first pre-train on a large unlabeled dataset, where effective feature representations can be learned. Examples of pre-training tasks include image inpainting, predicting rotations, and colorizing images. Then the final layers of the models can be optimized in specific downstream tasks, reducing the need for extensively labeled data. SimCLR  [11, 12] presents a seminal contrastive learning to train feature representation, where augmented views of an image (positive samples) are minimized within that image while being maximized against augmented views of other images (negative samples). MoCo  [13, 15, 14, 24, 25, 23] strengthens SSL with a dynamic dictionary, tapping into past batches for more negative samples, and a slowly updating momentum encoder for stable global representations. In contrast, Barlow Twins  [57] omits negative samples, focusing on refining features by measuring cross-correlation between positive samples. BYOL  [20] and DINO  [7] utilize a teacher-student setup, where the student refines its understanding by predicting the teacher’s representation, while the teacher network is updated through the slow-moving average of student’s parameters. Clustering-based methods like Deep Clustering  [5] and SwAV  [6] define pseudo-labels, sorting images into clusters, yielding great performance on downstream tasks.
In histopathology image analysis, self-supervised learning can be used to provide the backbone feature extraction module for MIL and patch-level classification models. Mingu et al.  [32] benchmarked SSL methods on pathology datasets, showcasing consistent enhancements in histopathology tasks. CS-CO  [56] also utilizes separate encoders for Hematoxylin images and Eosin images. Instead of contrastive loss between Hematoxylin images and Eosin images introduced in our work, CS-CO calculates a contrastive loss between different augmentations. In addition, CS-CO is a self-supervised learning method that requires encoders to generate a visual representation containing enough information to recover images in cross-stain prediction, while our semi-supervised CLASS-M only pursues shared latent features containing enough information to perform classification tasks. While self-supervised learning benefits from unlabeled data, semi-supervised learning has the advantage that it allows both unlabeled data and labeled data to be trained at the same time and use the knowledge learned from labeled data to better utilize unlabeled data.
Figure 1: The block diagram of our Contrastive Learning with Adaptive Stain Separation and MixUp (CLASS-M) for semi-supervised histopathological image classification. Orange part shows adaptive stain separation, where OD means Optical Density space. Green part shows mixup on both labeled and unlabeled samples.
@@ 2.3 Semi-supervised learning
Semi-supervised learning  [10, 43, 51] involves training a model simultaneously with both labeled and unlabeled data to enhance training outcomes, which reduces reliance on a limited number of annotated samples. Semi-supervised learning finds applications when obtaining fully labeled data is prohibitively expensive or impractical. For digital histopathological image analysis, getting a large amount of labeled data requires great effort from well-trained experts, which is expensive and time-consuming. However, it is much easier to acquire unlabeled data such as unannotated WSIs. One type of semi-supervised learning is consistency regularization which seeks agreement among the model predictions for the same input with different views, augmentations or epochs. For example, temporal ensembling  [34] aims to reach a consensus in the prediction of labels between current epochs and previous epochs.  [47] tries to make classifier’s prediction consistent across various transformations. More recent consistency regularization methods, such as FixMatch  [50] and MixMatch  [3], have been proposed. FixMatch assigns pseudo-labels to unlabeled samples when the prediction confidence is high with weak augmentation of input data. Then these pseudo-labels are used for training with strong augmentation. MixMatch is another consistency regularization method where pairs of labeled/unlabeled input data are linearly combined to create a "mixup" of data. The label of newly created data is a weighted average of the original labels followed by a sharpening process. Inspired by self-supervised learning, contrastive learning has also been integrated in semi-supervised learning  [1, 8, 49, 58]. Instead of seeking agreements in label predictions, contrastive learning focuses on the representation of positive and negative pairs. Positive pairs originate from the same inputs with different views or augmentations, while negative pairs involve different inputs. The method encourages a shorter distance between positive pairs and a longer distance between negative pairs in the feature representation space. In this paper, we propose a novel histopathology-specific contrastive learning based semi-supervised classification model.
@@ 3 Methods
3.1 An overview of CLASS-M
Figure 1 shows the workflow of CLASS-M. As shown in the orange part of Figure 1, slide-level stain separation is applied for generating Hematoxylin images and Eosin images from original RGB images. The original RGB images, Hematoxylin images and Eosin images are all augmented during training to improve the robustness of CLASS-M model. The model takes Hematoxylin images and Eosin images as input pairs to form different views of input data. The purple and pink boxes in Figure 1 illustrate the H and E ResNet  [26] encoders that have the same architecture but separate parameters to generate latent features $f_{H}$ and $f_{E}$. A contrastive loss is proposed for shared latent feature space between H channel and E channel to increase the similarity between $f_{H}$ and $f_{E}$. We take the average of features $f_{H}$ and $f_{E}$ and pass it to a linear+softmax layer to predict the labels. For labeled samples in training, the cross-entropy loss is introduced to measure the difference between predictions and true labels. Moreover, a challenge in semi-supervised learning in general is the limited number of labeled samples. Pseudo-labeling on unlabeled samples has been widely used in many state-of-the-art semi-supervised learning methods  [3, 50, 59]. In MixMatch  [3], after pseudo-labeling on original samples, the model further uses MixUp  [60] to introduce virtual samples by linear interpolation of two random samples. Inspired by those ideas and to fully utilize unlabeled samples, we first provide pseudo-labels to unlabeled samples, then add MixUp on both labeled and unlabeled samples to create virtual samples. The labels of the generated samples after MixUp are set to the weighted averages of original labels/pseudo-labels with a sharpening process. The green part in Figure 1 shows the workflow of this process for pseudo-labeling. We introduce different parts of CLASS-M model in detail in the following subsections.